# Bayesian mitigation of measurement errors in multi-qubit experiments

### v1

@@ II Theoretical background
Figure 3: (a): Quantum circuit for the preparation of a random multiqubit string: single qubit X gates are applied on random qubits. (b): Comparison between the mitigated (solid orange) and unmitigated (dashed gray) success probability of measuring the prepared bitstring as a function of the string length. (c): Mitigated (solid orange) success probability of measuring a prepared 16 qubits bitstring as a function of the number of shots. The results shown in (b) and (c) are averaged over 20 realizations of random bitstrings.
Figure 4: (a): Success probability of measuring a single qubit as a function of $\mathrm{N_{\mathrm{bin}}}$, representing the number of partitions used to construct the detector response function defined on the Q-axis. (b): Success probability of measuring a 20-qubit string as a function of $\mathrm{N_{\mathrm{bin}}}$. (c): Comparison of mitigated success probabilities for measuring a prepared bitstring as a function of string length, for different number of partitions of the Q-axis established when reconstructing the detector response functions. The results shown in (a), (b), and (c) are obtained with $10^{3}$ shots and averaged over 50 realizations.
@@ III Bayesian mitigation with binary outcomes
In this section, we apply our Bayesian measurement error mitigation strategy to binary multiqubit outcomes obtained by executing quantum circuits on the superconducting qubits quantum computer VTTQ20 (Leena, QPU from IQM). The protocol begins with single-qubit measurement tomography to determine the uncorrelated error rates. While the service provider typically provides these error rates, we prefer extracting them explicitly in our work. We characterize the detector by performing two simple experiments: qubit reset followed by measurement and qubit flip to the excited state followed by measurement. Using $\mathrm{N}_{\mathrm{shots}}=10^{5}$ shots for each experiment, we obtain a two-by-two noise matrix $\Lambda^{(q)}$ for each qubit in the QPU. With these, we have access to the multiqubit noise matrix and we are now able to apply the Bayesian measurement error mitigation algorithm to the outcome of any quantum circuit.
To showcase the effectiveness of the measurement error mitigation scheme, we select a benchmark algorithm designed to minimize errors from faulty gates and decoherence. Specifically, we use a multiqubit string preparation algorithm involving a single layer of single-qubit X-gates. As illustrated in Fig. 3(a), for a chosen bit string (e.g., ”10..101”), we apply X-gates to specific qubits to flip their state from ”0” to ”1”. In these circuits, noisy readout is the primary source of errors, with potential imperfections in the rotation gate calibration accounted for within the noise matrix.
@@
In Fig. 3(b), we display the success probability averaged over the 20 realizations as a function of the multiqubit string length, comparing it with the unmitigated probability. By comparing the mitigated results with the original unmitigated estimates, we can quantitatively assess the effectiveness of our error mitigation approach across different qubit lengths. We observe that the unmitigated success probability decreases rapidly as the string length increases, highlighting the challenge posed by measurement errors in longer sequences. On the other hand, applying the Bayesian measurement error mitigation results in significantly higher success probabilities across all lengths. Notably, for the longest string considered (19 qubits), the success probability improves from approximately $26\%$ without mitigation to $92\%$ after mitigation with just $10^{3}$ shots.
@@
Furthermore, in the context of this string preparation algorithm, the unmitigated result, on average, tends to plateau near the product of the individual readout fidelities, which for the 16 qubits used in this experiment corresponds to approximately $\sim 1/3$. In contrast, the Bayesian measurement error mitigation algorithm continuously refines with each shot, offering further advantages in its application. Although the rate of improvement diminishes over time, this behaviour is expected as Bayesian inference benefits from a larger number of measurements, allowing for a more precise estimation of the multiqubit state populations.
@@ IV Bayesian mitigation with Analog data
In this section, we demonstrate the enhancements to the Bayesian measurement error mitigation scheme that can be achieved by leveraging readout analog data from the detector. Here, to validate these improvements, we run the algorithms on the superconducting qubits quantum computer ”Nazca” - IBM127.
@@
where $\Lambda_{j}(Q)$ represents the probability of the state being $j$ when the detector measures the physical value $Q$, and $\mathbb{I}{[Q_{i},Q_{i+1}]}(Q)$ is the indicator function defined as $\mathbb{I}{[Q_{i},Q_{i+1}]}(Q)=1$ if $Q\in[Q_{i},Q_{i+1}]$ and 00 otherwise. $\lambda_{ji}$ is then the probability of the state being $j$ when the measured $Q$ is in the i-th interval $\left[Q_{i},Q_{i+1}\right]$. Here, $\mathrm{N_{\mathrm{bin}}}$ denotes the number of intervals used to discretize the Q-axis of the IQ-plane and serves as a parameter representing the discretization applied in modelling the detector response functions. While an analytical expression for these distributions would be ideal, and although a bimodal Gaussian can often adequately capture the behaviour of the readout functions, it is typically more straightforward and more efficient to use directly the histograms. This approach is somewhat similar to defining Q-dependent error rates along the Q-axis partition.
@@
Therefore, once we gather data from $\mathrm{N_{\mathrm{shots}}}=10^{5}$ shots, we establish a number of partitions, denoted as ${\mathrm{N_{\mathrm{bin}}}}$, and reconstruct response functions for each individual qubit. These response functions are then utilized in the multiqubit conditional probability of Eq. (8) to be used in the the update cycle of the Bayesian measurement error mitigation algorithm. In this framework, when we consider ${\mathrm{N_{\mathrm{bin}}}}=2$, this approach is equivalent to working with binary outcomes.
In Fig. 4, we investigate the influence of varying the number of partitions, ${\mathrm{N_{\mathrm{bin}}}}$, on Bayesian measurement error mitigation when using analog data. Therefore, we prepare 50 random bit strings across various qubit lengths and record the measurement outcomes in analog form. After obtaining the unmitigated initial estimates, we apply the Bayesian measurement error mitigation procedure, leveraging the previously characterized Q-dependent detector response functions. By examining how different values of ${\mathrm{N_{\mathrm{bin}}}}$ impact the mitigation process, we aim to optimize the partitioning strategy for improved readout error mitigation. This analysis allows us to understand the trade-offs between the granularity in the analog data and the effectiveness of the error mitigation, providing insights into how to best utilize the analog information obtained in quantum measurement. In Fig. 4(a), we present the mitigated single-qubit success probability averaged over the 50 realisation as a function of ${\mathrm{N_{\mathrm{bin}}}}$. We observe a noticeable trend where the success probability, which in the single qubit case is akin to the readout fidelity, improves as ${\mathrm{N_{\mathrm{bin}}}}$ increases. This trend becomes more pronounced and stable when analyzing longer bit strings. In Fig. 4(b), we display the mitigated success probability averaged over the 50 realisations for a 20-qubit string. Also in this case, we observe that the mitigated success probability increases as the response functions are discretized with a greater number of partitions ${\mathrm{N_{\mathrm{bin}}}}$. However, these improvements eventually plateau beyond a certain threshold value of ${\mathrm{N_{\mathrm{bin}}}}$.
In Fig. 4(c), we display the averaged mitigated success probabilities as a function of the string length for different values of ${\mathrm{N_{\mathrm{bin}}}}$. We see that increasing the granularity of the detector response function constantly result in better mitigated values for all string lengths. For the case of $\mathrm{N}_{q}=20$, the success probability significantly improves from approximately $\sim 89\%$ to about $\sim 94\%$ when the discretization of the response functions is increased from ${\mathrm{N_{\mathrm{bin}}}}=2$ to ${\mathrm{N_{\mathrm{bin}}}}=10$. Both cases represent a substantial improvement over the unmitigated result, which even though is not shown, it is around $\sim 25\%$. These findings underscore how leveraging the additional information encoded in the analog data can enhance the effectiveness of the Bayesian measurement error mitigation algorithm with minimal to none additional computational cost. Therefore, by optimizing the partitioning strategy for the detector analog data, we can achieve higher accuracy in the quantum measurements, thereby improving the overall performance of quantum algorithms on NISQ devices.
Figure 5: Comparison between success probabilities of measuring the prepared bit-strings as a function of the string length mitigated according to different readout error mitigation strategies. In solid blue and solid green we report the success probability mitigated trough Iterative Bayesian Unfolding and noise Matrix Invervesion Method respectively. In dashed orange and dashed yellow, we display the success probabilities mitigated trough our implementation of Bayesian measurement error mitigation using two partitions of the Q-axis for the response functions, $\mathrm{N_{bin}}=2$ and $\mathrm{N_{bin}}=20$ respectively. All the results displayed are obtained by averaging over 50 realisations.
V Comparison with other mitigation techniques
The results presented in the previous section provide compelling evidence of the efficacy of our implementation of Bayesian measurement error mitigation in reducing readout errors in multiqubit algorithms. However, it is relevant to compare the performances of our Bayesian approach with other established measurement error mitigation techniques. For benchmark purposes, we selected two techniques. The first benchmark measurement error mitigation scheme is the Iterative Bayesian Unfolding method (IBU) introduced at the end of Sec. II, which shares a probabilistic formulation linking correct qubit populations with the measured outcomes. However, IBU employs iterative expectation maximization rather than a formal prior-posterior update iteration in the Bayesian sense. The second measurement error mitigation technique is based on the Noise Matrix Inversion Method (MIM). This method applies the inverse of the noise matrix to the measured outcomes and, if the resulting mitigated probability vector is non-physical, minimize the squared distance squared $|\Lambda^{-1}\rho^{\mathrm{noise}}-\rho|^{2}$, enforcing the constraints of positivity and normalization.
In Fig. 5, we compare the success probabilities of the previously prepared random bitstrings mitigated using IBU, MIM, and our Bayesian measurement error mitigation method as a function of the multiqubit string length, averaged over 50 realizations. All methods effectively mitigate measurement errors, demonstrating their viability. For our Bayesian method we employ two extreme values for the discretization parameter $\mathrm{N}_{\mathrm{bin}}$ of the detector response function.
@@ V Comparison with other mitigation techniques
Figure 6: (a): Total variation distance between the mitigated probabilities after each Bayesian iteration as a function of the iteration number for different string lengths. (b): Time required by our Bayesian mitigation scheme to mitigate noisy outcomes as a function of the number of qubits measured. (c): Time required by our Bayesian mitigation scheme to mitigate the outcomes of the string preparation experiment run on 20 qubits as a function of the number of shots. The results are averaged over 50 realizations and correspond to the algorithm using $\mathrm{N}_{\mathrm{bin}}=20$ for the discretization of the detector response function.
When applying our Bayesian mitigation algorithm to binary outcomes, i.e fixing $\mathrm{N_{bin}}=2$, we achieve results very similar to IBU, with marginal improvements observed for the longer strings considered in our study. This similarity is expected because both methods start from similar probabilistic foundations but employ different iterative approaches to reach the optimization/maximization. On the other hand, by increasing the discretization parameter of the response function, our Bayesian method outperforms IBU. This enhancement suggests that our approach effectively leverages additional information encoded in the readout signal, which contributes to improved mitigated outcomes. Comparing our Bayesian method with MIM reveals a peculiar trend. Initially, MIM consistently outperforms our Bayesian method when applied to binary outcomes. However, with $\mathrm{N_{bin}}=20$, we observe a transition across different qubit numbers: MIM excels for shorter strings, whereas our method begins to surpass MIM for longer strings. Specifically, this transition appears for qubit numbers greater than 5 in our experimental results.
These comparisons highlight the strengths of each method under different conditions, showcasing how our Bayesian measurement error mitigation approach applied to the detector analog data can adapt and potentially outperform more established traditional methods like IBU and MIM when mitigating the outcome of large multiqubit experiments.
VI Scaling analysis and other considerations
@@
To conclude, we estimate the computational complexity for the Bayesian multiqubit measurement error mitigation algorithm 2. A single Bayesian iteration, as described in Eq. (2), scales linearly with the number of shots $\mathrm{N}_{\mathrm{shots}}$. This process must be repeated for each pair of states in the truncated subspace. If initially we have $M$ strings with counts such that $\rho_{\bm{i}}^{\mathrm{noisy}}\neq 0$, we have $\frac{M(M-1)}{2}$ populations pairs to consider and update. Thus, if all pairs are updated $N_{r}$ times, the overall computational complexity to post-process the results of a multiqubit experiment scales as $\propto\frac{N_{r}\mathrm{N}_{\mathrm{shots}}M(M-1)}{2}$. This indicates a linear relationship with the number of shots and iterations and a quadratic relationship with the number of states after truncation based on the initial noisy counts. Therefore, the primary computational cost for large systems arises from the number of initial noisy populations, which is intricately linked to the number of shots and the extent to which the quantum state is spread over the multiqubit Hilbert space. Hence, the algorithm is more efficient when the outcome of the quantum algorithm is concentrated on a limited number of populations.
@@
To determine the necessary number of updates for the multiqubit state populations, Fig. 6 (a) shows the total variation distance between the probability vectors before and after each Bayesian update from Eq. (7) for each population pair. The results for qubit counts of $\mathrm{N}_{q}=5,10$ and 20 indicate that convergence is achieved after a few iterations as the figure of merit reaches a plateau. Consequently, we set a threshold to terminate the Bayesian measurement error mitigation algorithm when the total variation distance between probability vectors before and after an iteration falls below $\epsilon=10^{-3}$. To ensure the algorithm exits the iteration loop, we imposed a maximum of 20 iterations, which was never reached in our simulations.
@@
With this exit condition, we recorded the time required to run the Bayesian mitigation algorithm. In Fig. 6 (b) and (c), we display the average time to mitigate the outcome of the random string state preparation algorithm as a function of the number of qubits and the number of shots for a circuit with 20 qubits, respectively. The timing includes all the steps necessary to perform the mitigation algorithm as detailed in 2. Using a conventional quad-core laptop, the post-processing of the experiment data was within 1 second in the worst case considered here, i.e. $\mathrm{N}_{q}=20$ and $\mathrm{N}_{\mathrm{shots}}=2000$.
VII Conclusions

### v2

@@ II Theoretical background
Figure 3: (a): Quantum circuit for the preparation of a random multiqubit string: single qubit X gates are applied on the specific qubits ti flip. (b): Comparison between the mitigated (orange) and unmitigated (gray) success probability of measuring the prepared bitstring as a function of the string length. (c): Comparison between the mitigated (orange) and unmitigated (gray) success probability of measuring a prepared 16 qubits bitstring as a function of the number of shots. The distribution shown in (b) and (c) correspond to 20 different random bitstrings, with the middle bar representing the mean across the different realization, while the upper and lower bars are the highest and lowest success probabilities.
Figure 4: (a): Comparison of mitigated success probabilities for measuring a prepared bitstring as a function of string length, for different number of partitions of the Q-axis. The results are obtained with $10^{3}$ shots and averaged over 50 realizations, with error bars corresponding to the standard deviations. (b): Success probability of measuring a single qubit as a function of $\mathrm{N_{\mathrm{bin}}}$ with the inset showing only the average values. (c): Success probability of measuring a 20-qubit string as a function of $\mathrm{N_{\mathrm{bin}}}$. The results shown in (b) and (c) are obtained with $10^{3}$ shots and show the results for all the 50 realizations. The middle horizontal bar represent the average mitigated success probability, while the upper and lower bars are the maximum and minimum of the measured samples.
@@ III Bayesian mitigation with binary outcomes
In this section, we apply our Bayesian measurement error mitigation strategy to binary multiqubit outcomes obtained by executing quantum circuits on the superconducting qubits quantum computer VTTQ20. The protocol begins with single-qubit measurement tomography to determine the uncorrelated error rates. While the service provider typically provides these error rates, we prefer extracting them explicitly. We characterize the detector by performing two simple experiments: qubit reset followed by measurement and qubit flip followed by measurement. Using $\mathrm{N}_{\mathrm{shots}}=10^{5}$ shots for each experiment, we obtain a two-by-two noise matrix $\Lambda^{(q)}$ for each qubit in the QPU. With these, we have access to the multiqubit noise matrix and we are now able to apply the Bayesian measurement error mitigation algorithm to the outcome of any quantum circuit.
To showcase the effectiveness of the measurement error mitigation scheme, we select a benchmark algorithm designed to minimize errors from faulty gates and decoherence. Specifically, we prepare a specific bit string through a single layer of single-qubit X-gates, i.e $\mathrm{U}(\mathbf{s})\left|{\mathbf{0}}\right\rangle=\otimes_{i}X^{s_{i}}_{i}\left|{\mathbf{0}}\right\rangle$. As illustrated in Fig. 3 (a), for a chosen bit string (e.g., ”10..101”), we apply X-gates to specific qubits to flip their state from ”0” to ”1”. The benchmark figure of merit is the success probability, which is the probability of measuring the specific string $\mathbf{s}$ after applying $\mathrm{U}(\mathbf{s})$ to the initial state. In these circuits, noisy readout is the primary source of errors, with potential imperfections in the rotation gate calibration accounted for within the noise matrix.
@@
In Fig. 3 (b), we display the success probabilities obtained over 20 realizations as a function of the multiqubit string length, comparing them against the unmitigated ones. We observe that the unmitigated success probabilities decrease rapidly as the string length increases, highlighting the challenge posed by measurement errors in longer sequences. However, applying Bayesian measurement error mitigation results in significantly higher success probabilities across all lengths. In particular, for the longest string considered (19 qubits), the average success probability improves from approximately $26\%$ without mitigation to $92\%$ after mitigation with only $10^{3}$ shots. Furthermore, we observe that mitigated values have a narrower distribution around the mean for each qubit number.
@@
In the context of this string preparation circuit, the unmitigated result, on average, tends to plateau near the product of the individual readout fidelities, which for the 16 qubits used in this experiment corresponds to approximately $\sim 1/3$. In contrast, the Bayesian measurement error mitigation algorithm continuously refines the mitigated outcome with each shot, offering further advantages in its application. Although the rate of improvement diminishes over time, this behavior is expected as Bayesian inference benefits from a larger number of measurements, allowing for a more precise estimation of the multiqubit state populations.
@@ IV Bayesian mitigation with Analog data
In this section, we demonstrate how readout analog data from the detector can boost the accuracy of the Bayesian measurement error mitigation scheme. Here, to validate these improvements, we run the algorithms on the superconducting qubits quantum computer ”Nazca” - IBM127 (for details on calibration and QPU figures of merit see the Supplemental Material [45]).
@@
where $\Lambda_{j}(Q)$ represents the probability of the state being $j$ when the detector measures the physical value $Q$, and $\mathbb{I}{[Q_{i},Q_{i+1}]}(Q)$ is the indicator function defined as $\mathbb{I}{[Q_{i},Q_{i+1}]}(Q)=1$ if $Q\in[Q_{i},Q_{i+1}]$ and 0 otherwise. $\lambda_{ji}$ is then the probability of the state being $j$ when the measured $Q$ is in the i-th interval $\left[Q_{i},Q_{i+1}\right]$. Here, $\mathrm{N_{\mathrm{bin}}}$ denotes the number of intervals used to discretize the Q axis of the IQ plane and serves as a parameter representing the discretization applied in the modeling of the detector response functions. Although an analytical expression for these distributions would be ideal, and although a bimodal Gaussian can often adequately capture the behavior of the readout functions, it is typically simpler to use directly the histograms. This approach is somewhat similar to defining Q-dependent error rates along the Q-axis partition.
@@
Therefore, once we collect data from the $\mathrm{N_{\mathrm{shots}}}=10^{5}$ shots, we establish a number of partitions, denoted as $\mathrm{N_{\mathrm{bin}}}$}, and reconstruct response functions for each individual qubit. These response functions are then used in the multiqubit conditional probability of Eq. (8) to be used in the the update cycle of the Bayesian measurement error mitigation algorithm. In this framework, when we consider ${\mathrm{N_{\mathrm{bin}}}}=2$, this approach is equivalent to working with binary outcomes.
@@
Figure 5: (a): Mitigated success probability as a function of the Hamming distance between population pairs undergoing the Bayesian update cycle. (b): Total variation distance between the mitigated probabilities after each Bayesian iteration as a function of the iteration number for different string lengths. (b): Number of iterations required by our Bayesian mitigation scheme to reach the convergence condition. (c): Time required by our Bayesian mitigation scheme to mitigate the outcomes of the string preparation experiment as a function of the number of qubits measured. The results in (a), (b), and (c) show the results obtained for 50 realizations with the middle horizontal bar line representing the mean of the distributions. In all cases the mitigation algorithm is applied using $\mathrm{N}_{\mathrm{bin}}=20$ as discretization parameter for the detector response function.
In Fig. 4, we investigate the role of the number of partitions, $\mathrm{N_{\mathrm{bin}}}$}, on Bayesian measurement error mitigation when using analog data. We prepare 50 random bit strings across various qubit lengths and record the measurement outcomes in analog form. After obtaining the unmitigated initial estimates, we apply the Bayesian measurement error mitigation procedure, leveraging the previously characterized Q-dependent detector response functions. Our aim is to understand how different values of $\mathrm{N_{\mathrm{bin}}}$}, i.e. the granularity of the analog measurement data, can impact the mitigation process. In Fig. 4 (a), we display the averaged mitigated success probabilities as a function of the string length for different values of $\mathrm{N_{\mathrm{bin}}}$} with the respective standard deviations. We see that increasing the granularity of the detector response function constantly result in better mitigated values for all string lengths. For the case of $\mathrm{N}_{q}=20$, the success probability improves significantly from approximately $\sim 89\%$ to about $\sim 94\%$ when the discretization of the response functions increases from ${\mathrm{N_{\mathrm{bin}}}}=2$ to ${\mathrm{N_{\mathrm{bin}}}}=10$. Both cases represent a substantial improvement over the unmitigated result which is around $\sim 25\%$. In Fig. 4(b), we show the mitigated single-qubit success probability for all the the 50 realisations as a function of $\mathrm{N_{\mathrm{bin}}}$}. In the inset, which displays only the average values, we observe a noticeable trend where the success probability, which in the single qubit case is akin to the readout fidelity, improves as $\mathrm{N_{\mathrm{bin}}}$} increases. In Fig. 4(c), we display the mitigated success probability for all the 50 realisations for a 20-qubit string. Also in this case, we observe that the mitigated success probability increases as the response functions are discretized with a greater number of partitions $\mathrm{N_{\mathrm{bin}}}$}. However, these improvements eventually plateau beyond a certain threshold value of $\mathrm{N_{\mathrm{bin}}}$}. These findings underscore how leveraging the additional information encoded in the analog data can enhance the effectiveness of the Bayesian measurement error mitigation algorithm with minimal to none additional computational cost.
IV.1 Algorithm convergence and scaling
@@
In this subsection, we estimate the computational complexity and cost of the Bayesian multiqubit measurement error mitigation algorithm II and applications of the previous subsection. A single Bayesian iteration, as described in Eq. (2), scales linearly with the number of shots $\mathrm{N}_{\mathrm{shots}}$. This process must be repeated for each pair of states in the truncated subspace. If initially we have $M$ strings with counts such that $\rho_{\bm{i}}^{\mathrm{noisy}}\neq 0$, we have $\frac{M(M-1)}{2}$ populations pairs to consider and update. Thus, if all pairs are updated $N_{r}$ times, the overall computational complexity to post-process the results of a multiqubit experiment scales as $\propto\frac{N_{r}\mathrm{N}_{\mathrm{shots}}M(M-1)}{2}$. This indicates a linear relationship with the number of shots and iterations and a quadratic relationship with the number of states after truncation based on the initial noisy counts. Therefore, the primary computational cost for large systems arises from the number of initial noisy populations, which is intricately linked to the number of shots and the extent to which the quantum state is spread over the multiqubit Hilbert space. Hence, the algorithm is more efficient when the outcome of the quantum algorithm is concentrated on a limited number of populations. Thus, it is convenient to introduce another restriction and send through the Bayesian update cycle only pairs whose bit strings are within a certain Hamming distance $d$, e.g. $H_{d}(i,j)\leq d$. In Fig. 5 (a), we display the mitigated success probability for 20 qubits strings as a function of the Hamming distance used for selecting the qubit pairs. We see how, on average, there is negligible improvement after a Hamming distance of 3, which is then the threshold we set in all the executions of the Bayesian mitigation algorithm if not stated otherwise.
@@
Fig. 5 (b) shows the total variation distance between the probability vectors before and after each Bayesian update from Eq. (7) for each population pair. The results shown here for a single realisation for qubit counts of $\mathrm{N}_{q}=20,40$ and 60 indicate that convergence is achieved after a few iterations as the figure of merit goes below $10^{-17}$. Consequently, we set a threshold to terminate the Bayesian measurement error mitigation algorithm when the total variation distance between probability vectors before and after an iteration falls below $\epsilon=10^{-8}$. To ensure the algorithm exits the iteration loop, we imposed a maximum of 20 iterations, which was never reached in our simulations.
With this exit condition, we recorded the number of iteration required for our Bayesian mitigation algorithm to converge. In Fig. 5 (c), we show the number of iterations required for all 50 realizations and notice how the algorithm reaches the thresholds within 7 steps for all the case considered here. On average, it seems that the number of iterations slowly increases with the number of qubits.
@@
In Fig. 5 (c), we display the corresponding time required to mitigate the outcome of the random string state preparation circuit as a function of the number of qubits. The timing includes all the steps necessary to perform the mitigation algorithm as detailed in II. Using a conventional quad-core laptop, the post-processing of the experiment data was on average within 5 second in the worst case considered here, i.e. $\mathrm{N}_{q}=60$ and $\mathrm{N}_{\mathrm{shots}}=1000$.
Figure 6: (a): Comparison between success probabilities of measuring the prepared bit-strings as a function of the string length mitigated according to different readout error mitigation strategies. The results displayed are obtained by averaging over 50 realisations with error bars corresponding to the standard deviations. (b): Distribution over the 50 realisations for the mitigated success probabilities the different mitigation algorithms. (c) Comparison between the unmitigated and mitigated GHZ state fidelity calculated according through multiqubit coherence as a function of the number of qubit and mitigated through different mitigation strategies. The results shown in (b) and (c) are obtained with $10^{3}$ shots. The data shown in (a) and (b) was obtained on IBM127, while circuits for (c) were executed on VTTQ50.
V Comparison and Integration with other mitigation techniques
The previous sections provide evidence about the performances of our implementation of Bayesian measurement error mitigation when applied to multiqubit experiments. However, it is important to compare our method with other established measurement error mitigation techniques and for other quantum algorithms. For benchmark purposes, we selected two techniques. The first benchmark measurement error mitigation scheme is the Iterative Bayesian Unfolding algorithm (IBU) introduced at the end of Sec. II, which shares a probabilistic formulation with our protocol, linking correct qubit populations with the measured outcomes. However, IBU employs iterative expectation maximization rather than a formal prior-posterior update iteration in the Bayesian sense. The second measurement error mitigation technique is the Mthree method [37], based on noise matrix inversion. This method applies the inverse of the noise matrix to the measured outcomes in a reduced sub-space and, if the resulting mitigated probability vector is non-physical, finds the closest positive vector which still satisfies the sum-to-one condition. In Fig. 6 (a), we compare the success probabilities of measuring the prepared random bitstrings mitigated using IBU, Mthree, and our Bayesian measurement error mitigation method (dubbed in the figure as BMEM) as a function of the multiqubit string length, averaged over 50 realisations. All methods mitigate measurement errors efficiently, demonstrating their validity. We apply three version of our Bayesian method: two cases where we apply our mitigation algorithm using the noisy counts as initial population pairs and two values for the discretization parameter $\mathrm{N}_{\mathrm{bin}}$ of the detector response function, i.e. 2 and 20; one case where the Bayesian mitigation algorithm uses the Mthree mitigated values as initial estimates for the multiqubit populations and $\mathrm{N}_{\mathrm{bin}}$=20. The workflow of the latter is depicted in Fig. 1 (c).
When applying our Bayesian mitigation algorithm to binary outcomes, i.e fixing $\mathrm{N_{bin}}=2$, we achieve results very similar to IBU, with marginal improvements observed for the longer strings considered in our simulations. This similarity is expected because both methods start from similar probabilistic foundations, but employ different iterative approaches to reach the optimization/maximization. On the other hand, by increasing the discretization parameter of the response function, our Bayesian method outperforms IBU. This enhancement suggests that our approach effectively leverages additional information encoded in the readout signal, which contributes to improved mitigated outcomes. Furthermore, Mthree seems to perform better than IBU, but on average worse than our Bayesian approach using analog data. Moreover, the highest accuracy is obtained by combining our Bayesian scheme using analog data with the Mthree mitigated outcomes as initial estimates for the Bayesian update cycle. In Fig. 6 (b), we display the distributions of the mitigated success probabilities for the $\mathrm{N_{q}}=10$ qubits bitstrings where we observe how the different realisations are distributed around the average. It is worth to mention again how IBU and the Bayesian mitigation applied with binary outcomes have a very similar profile. As a next benchmark example, we prepare a GHZ state and calculate the state fidelity through the multiple quantum coherence [46, 47]. In Fig. 6 (c), we compare the unmitigated state fidelity with the ones obtained mitigating the circuits outcomes with the different mitigation techniques. All the techniques used show an improvement over the unmitigated result. As in the previous example, Mthree and Bayesian measurement mitigation with analog data seem to perform better than IBU, while Mthree results in the highest accuracy for most qubit numbers. To conclude our benchmark, in Fig. 7, we compare the time employed by the Bayesian measurement error mitigation algorithm when mitigating the outcome of three different quantum states as a function of the qubit number. The three states considered are: random bit-string, GHZ state, and uniform superposition, i.e. the state obtained by applying an Hadamard gate to each qubit. We observe how the computational time is lower when mitigating single multiqubit strings, and increases when mitigating the outcome of a GHZ state and uniform superposition. The latter can be considered an upper bound, or worst case scenario, to the computational cost of the algorithm, as the solution is spread over the entire basis state, and the number of different basis states included in the mitigation algorithm is determined by the number of shots.
Figure 7: Time required by our Bayesian mitigation scheme to mitigate the outcomes for different quantum states as a function of the number of qubits. Solid lines consider the case where the Bayesian mitigation algorithm is applied starting from the noisy counts, the dashed to case where it is applied starting from the Mthree estimates as initial counts. All three state preparations were executed on VTTQ50 and $10^{3}$ shots.
VI Conclusions